# Contact SLAM: An Active Tactile Exploration Policy Based on Physical Reasoning Utilized in Robotic Fine Blind Manipulation Tasks

### v1

@@ V-C Results of the Socket Assembly Experiments
The experimental results are presented in Fig. 6. As can be observed, by leveraging variations in tactile signals and the proposed active exploration framework, the robot is able to achieve incremental recognition of relative contact regions. Through interaction, the robot progressively reduces uncertainty in its understanding of the scene. Ultimately, after $6\sim 8$ exploration steps, the estimated relative contact distribution between the grasped object and the environment converges from a multi-modal distribution to a uni-modal distribution, thereby enabling the construction of a closed-loop control scheme to successfully accomplish the peg-in-hole assembly task.
@@
After the localization process, the robot analyzed the pose errors. then moved the plug to the target region, and finished the assembly task with a spiral hole search strategy finally.
We futher test the accuracy and efficiency of the proposed contact SLAM method in the above task scenarios. We made each task 5 times, and the results are in Table I and Fig. 7.
TABLE I: The results of socket assembly experiments.
Mean Errors(mm) Mean Iteration Number
two-pin $3.775$ $7.13$
three-pin $1.815$ $7.67$
Figure 7: The Uncertainties of the scenarios during the socket assembly task. (a) represents the two-pin plug experiment, while (b) represents the three-pin plug experiment. Distribution std means the standard deviation of the particles’ distribution. Particle count means the number of remaining particles at each iteration step.
It could be found in the Table. I that after 6 to 8 iterations, the localization error remained 1 to 5mm, which is related to the error of resampling of particles, as well as the localization error between the gripper and the grasped plug. When re-sampling new particles, in order to cover the accurate pose of the object, the distribution of particles is random in a certain range, causing potential errors. Whatever, during the assembly process, the localization error was reduced gradually, as illustrated in Fig. 7, the Distribution of particles reduced from 30mm to within 5mm.
V-D Results of the Blind Pushing Experiments
@@
As shown in Fig. 8, in this task scenario, the robot is capable of detecting the relative contact region between the T-shaped tool and the block using the method described in Section V-B, which enables continuous monitoring of the block’s position throughout the pushing process. When obstacles are encountered, the robot is able to reduce the localization error of the obstacles to within 10 mm through a limited number of active explorations and contact interactions, which is sufficient to meet the requirements of trajectory re-planning.
V-E Experimental Results Discussion

### v2

@@ IV-B Results of the Socket Assembly Experiments
IV-B1 The ablation experiments
We introduce two improvements to the particle filtering procedure, namely a distance variance criterion and adaptive resampling. The necessity of these improvements is demonstrated through ablation experiments. Specifically, for each type of socket pin, we select a random initial pose, and conduct the control group(fixed resampling without the distance variance criterion) and two experimental groups, one using the distance criterion alone and the other combining the distance criterion with adaptive resampling.
For the control group, since entropy is the only information criterion, when the particle distribution becomes highly concentrated, the resulting motion direction is unique in all cases, as shown in Fig. 5. This indicates that the entropy criterion fails under such conditions. To address this issue, we introduce the distance variance criterion as shown in Section III-D. The specific implementation is as follows: when all entropy values fall below $10^{-2}$, the decision criterion is switched from entropy to the distance criterion. We then reattempt exploration in the previously failed scenarios. As shown in Fig. 5, the results demonstrate that, after introducing the distance criterion, effective environment exploration can still be achieved even when the entropy criterion is extremely small, thereby leading to correct localization results.
Figure 5: The comparison of the control group and experimental group adding the distance criterion. In iteration 4, the entropy criterion is small and unique in all directions, indicating that the active exploration failed. After introducing the distance criterion, the optimal information gain is obtained.
In the control group, a resampling strategy was adopted in which resampling is triggered only when the number of particles falls below 50. Under this strategy, accumulated particle position errors in earlier stages may result in certain particle modes being eliminated from the new particle set before the resampling threshold is reached, as illustrated in Fig. 6, ultimately causing localization failure. We added the adaptive resampling to the filter process as shown in Section III-D, and the rule of resampling is as follows:
$scale$ $=$ $\left\{\begin{matrix}\frac{P_{t}}{3\times P_{t+1}}&len(P_{t+1})>300\\ \frac{P_{t}}{2\times P_{t+1}}&100\leq len(P_{t+1})\leq 300\\ \frac{P_{t}}{P_{t+1}}&50\leq len(P_{t+1})<100\\ \frac{2\times P_{t}}{P_{t+1}}&len(P_{t+1})<50\end{matrix}\right$. (10)
$P_{t+1}^{\prime}=scale\times P_{t+1}$ (11)
The results are shown in Fig. 6. It can be observed that, after incorporating adaptive resampling, the particle distribution was better maintained, and the localization process successfully converged to a unimodal distribution.
Figure 6: The comparison of the control group and experimental group utilizes adaptive resampling. In iteration 5, the particle set around the plug was eliminated because of the accumulated position errors during exploration. In the experimental group,
IV-B2 The comparative experiments
To verify the accuracy of the proposed method for estimating the relative position of the peg and the hole, we conducted 25 sets of experiments for each pin assembly task. The initial position of each set of experiments was randomly selected within the plane of the hole. In total, 50 sets of experiments were carried out for data analysis. For the control group, we also conducted 25 sets of experiments for each pin at arbitrary initial positions, making a total of 50 sets of experiments for comparison. For the experimental group, when the estimation of the relative position between the plug and the socket converges to a unimodal distribution, we use the method as described in the article [11] for the final assembly operation. For the spiral hole-searching method used for comparison, we set the following threshold conditions: $F_{x/y}>0.55*F_{z}$ if $F_{z}>0.5$; $F_{x/y}>0.9$ if $F_{z}<0.5$. If the threshold conditions are met, it is considered that the hole area has been found, and then the plug is inserted downward and it is checked whether the assembly is completed. For both the experimental group and the control group, the conditions for successful assembly are: $d_{z}>0.025$ and $F_{x/y}>0.9N$.
TABLE I: The results of socket assembly experiments.
RMSE Iter.Num Explor.time SR
two-pin $0.449mm$ $7.70$ $507.31s$ $80\%$
three-pin $0.698mm$ $9.95$ $679.96s$ $84\%$
sprial search $165.04s$ $10\%$
The experimental results are shown in Tab. I. It can be seen that the success rate of assembly using the contact positioning and active exploration tactile method is significantly higher than that of the control group. For the assembly tasks of the two types of bolts we selected, the success rates reached $80\%$ and $84\%$ respectively. In contrast, the spiral hole-searching method used for comparison only succeeded once in all attempts. This indicates that compared with the passive hole-searching method, the hole-searching method with active reasoning ability has significant advantages when facing an environment with redundancy and interference. The hole-searching method with active reasoning ability can model the environmental uncertainty, determine the possible potential positions, and then further narrow down the possible potential positions during continuous interaction, gradually eliminating the uncertainty. On the other hand, the passive search method is only suitable for situations with small errors. When there are large errors and environmental interference, due to the limitations of observation, it is likely to miss the correct position and fall into local optima.
Fig. 7 intuitively shows the process of relative position positioning using contact-based positioning and active tactile exploration. It can be seen that when the edge of the plug collides with the edge of the hole or goes beyond the hole area, the positions and weights of the particles are redistributed. Particles with too low weights or outside the potential distribution area are deleted. After $7\sim 9$ explorations, the distribution of the particles gradually stabilizes. If there is a single-peak distribution at this time, it indicates that the exploration is successful. Failed cases are mainly concentrated in situations where the active exploration fails to converge, that is, there is still a multi-peak distribution after the maximum number of exploration steps. The reason for the non - convergence of the exploration is that the hole area is complex and the sizes of different areas are similar. If the plug falls into a hole area with a similar size, the exploration movement range will be greatly limited. If the plug has not explored the overall environmental area comprehensively at this time, the particles will remain in the hole areas with similar sizes and cannot converge to a single-peak distribution.
@@ IV-B2 The comparative experiments
To verify the accuracy of the position localization results of contact SLAM, we verified the position estimation results of the socket obtained during the active exploration process. The results are shown in Fig. 8.
Figure 8: The estimation results of the socket. Different initial position of gripper and plug got the same pose estimation of socket after several iterations.
The reason why we chose to estimate the socket position in the environment as a verification of the positioning accuracy of contact SLAM is that during the task execution process, the positions of the gripper and the grasped object are constantly changing, and only the environment remains unchanged. As can be seen from Fig. 8, with the convergence of the particle distribution, the results of socket position estimation starting from different initial positions will converge to the vicinity of the true value after exploration, and the error is within 2 mm, indicating the effectiveness of our contact SLAM for unknown estimation.
IV-C Results of the Blind Pushing Experiments
@@
As shown in Fig. 9, in this task scenario, when obstacles are encountered, the robot is able to reduce the localization error of the obstacles to within 10 mm through a limited number of active explorations and contact interactions, which is sufficient to meet the requirements of trajectory re-planning.
IV-D Experimental Results Discussion